OpenStudy (anonymous):

A single tree produces about 2.6 × 102 lb of oxygen each year. The Amazon rainforest has about 3.9 × 1011 trees. About how many pounds of oxygen are produced in the rainforest each day ? 2.8 × 10 lb

OpenStudy (shreenithaa):

(3.9x1011)times(2.6x102) whats the answer?

OpenStudy (anonymous):

i have to type it in

OpenStudy (shreenithaa):

what do u mean?

OpenStudy (anonymous):

I have to type the answer myself

OpenStudy (shreenithaa):

in what

OpenStudy (shreenithaa):

anyways do you have the answer in multiplication form or as in one solution form

OpenStudy (anonymous):

can u help me get the answer plz and hurry I still have a lot of work to do im in homeschool

OpenStudy (shreenithaa):

k wait..

OpenStudy (anonymous):

102=10^2 1011=10^11

OpenStudy (shreenithaa):

1045657.08 pounds of oxygen

OpenStudy (shreenithaa):

what's 102=10^2 1011=10^11( this all about

OpenStudy (anonymous):

that wont all fit

OpenStudy (anonymous):

in the answer box

OpenStudy (shreenithaa):

can u do it in scientfic notation?

OpenStudy (anonymous):

do u know anybody in k12 that did this lesson already

OpenStudy (shreenithaa):

maybe @jennyrlz @zepdrix

OpenStudy (shreenithaa):

what grade r u in

OpenStudy (anonymous):

8th

OpenStudy (shreenithaa):

me too

OpenStudy (anonymous):

u in k12?

OpenStudy (shreenithaa):

what is that

OpenStudy (anonymous):

nevermind

OpenStudy (anonymous):

can u invite those other people to answer this question

OpenStudy (shreenithaa):

i just did by writing @ and their username

OpenStudy (shreenithaa):

the blue thingies mean i just asked them

OpenStudy (anonymous):

oh im new srry

OpenStudy (shreenithaa):

its k @Hero can u help

hero (hero):

@master479 are you taking a test?

OpenStudy (anonymous):

yes plz i need ur help

OpenStudy (shreenithaa):

check this out i think this may contain ur answer 54a59c97e4b0e4319f8f9904

OpenStudy (shreenithaa):

i mean http://openstudy.com/study#/updates/54a59c97e4b0e4319f8f9904

OpenStudy (anonymous):

is that a link? its not in blue

OpenStudy (shreenithaa):

no the other one

OpenStudy (anonymous):

I looked at that already I dont trust it

OpenStudy (shreenithaa):

http://brainly.com/question/116747 this one has the same answer

OpenStudy (anonymous):

but do I put 1.014 or the 14

OpenStudy (shreenithaa):

u put 1.014*10^{14}

OpenStudy (anonymous):

k ill tell u the % i got brb

OpenStudy (anonymous):

it wont even fit

OpenStudy (anonymous):

ill just put 14

OpenStudy (shreenithaa):

i dont know sorry

OpenStudy (anonymous):

80% I PASSED!!!!!!!!! YAYYYYYYYYYYY

OpenStudy (anonymous):

school is so stressful

OpenStudy (anonymous):

I medadled u

OpenStudy (shreenithaa):

yayy was i right?

OpenStudy (shreenithaa):

r u in homeschool

OpenStudy (anonymous):

sorry I suck at spelling I dont know I cant check without the password and I dont know it

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

im in homeschool

OpenStudy (shreenithaa):

yay.. wait is the yes for im right or..

OpenStudy (anonymous):

yes im in homeschool

OpenStudy (shreenithaa):

oh so i was wrong sorry

OpenStudy (anonymous):

u knew the answer!!!!!!!!!!!!!

OpenStudy (shreenithaa):

OMG BUT ACtUALLY the useful INTERNET did

OpenStudy (anonymous):

r u in homeschool

OpenStudy (shreenithaa):

nope

OpenStudy (anonymous):

k bye luv

OpenStudy (shreenithaa):

bye what state r u in

OpenStudy (anonymous):

sorry dont give to much personal information

OpenStudy (shreenithaa):

k tata..

OpenStudy (shreenithaa):

u can close ur question now

hero (hero):

posting test questions on OS results In automatic suspension